the value of a 2 digit number is more than 3 times the sum of its digits. The units digit is 1 more than twice the tens digit. find the 2 digit number.
Answer by ankor@dixie-net.com(12701)

I had to add some missing information here, hopefully it's correct
:
let 10x + y = the two digit number
:
Write an equation for each statement:
:
"the value of a 2 digit number is 1 more than 3 times the sum of its digits."
10x + y = 3(x + y) + 1
10x + y = 3x + 3y + 1
10x - 3x = 3y - y + 1
7x = 2y + 1
:
"The units digit is 1 more than twice the tens digit."
y = 2x + 1
:
find the 2 digit number.
Substitute (2x+1) for y
7x = 2(2x+1) + 1
7x = 4x + 2 + 1
7x - 4x = 3
3x = 3
x = 1
then
y = 2(1) + 1
y = 3
:
13 is the number
;
:
See if that's true in the statement:
"the value of a 2 digit number is 1 more than 3 times the sum of its digits."
13 = 3(1 + 3) + 1
13 = 3(4) + 1